OpenStudy (anonymous):

What are the solutions of x2 + 2x − 24 = 0?

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to -24 and add to 2 at the same time.

jimthompson5910 (jim_thompson5910):

What are those two numbers?

OpenStudy (anonymous):

i have no idea

jimthompson5910 (jim_thompson5910):

list out the factors of 24

OpenStudy (anonymous):

x^2-2x-24=0 (x-6)(x+4)=0 x-6=0 or x+4=0 x=6 or x=-4 <---- Answers